OpenStudy (liz_beauty45):

What is the LCM of 9 and 12? A) 3 B) 9 C) 18 D) 36

OpenStudy (mrm):

Have you tried using any formulas to solve this problem?

OpenStudy (liz_beauty45):

no but i think it may be B or C

OpenStudy (mrm):

Hmm... well lets talk about it first, we'll get to the bottom of it. Did you know that you can find the LCM of 2 numbers a and b by multiplying them together, then dividing by thir greatest common denominator (GCD)? You can also you a technique called prime factorization which is useful for finding the LCM of many numbers.

OpenStudy (mrm):

Let's use the first method. \[LCM = \frac{ ab }{ GCD(a,b) }\]

OpenStudy (liz_beauty45):

ohhh i didnt know

OpenStudy (liz_beauty45):

mhmm

OpenStudy (mrm):

if we let a= 9 and b = 12, what is GCD(a,b)?

OpenStudy (thebrainzzzzz):

I think D.

OpenStudy (thebrainzzzzz):

Because 9, 18, 27, (36), 45 and for 12 its 12, 24 (36), 48

OpenStudy (liz_beauty45):

um either 21 or..somethingelse NO NOT 21 CUZ its not an option

OpenStudy (liz_beauty45):

but maybe 36??

OpenStudy (thebrainzzzzz):

Yes

OpenStudy (liz_beauty45):

@mrm is that right?

OpenStudy (studydiva):

ITS 36 BECAUSE 9*4=36 AND 12*3=36

OpenStudy (thebrainzzzzz):

I just said that -_-

OpenStudy (liz_beauty45):

thankyou so much everyone yall are AMAZING PEOPLE or robots....but yeah thanks!!

OpenStudy (welshfella):

another way is to take the prime factors 9 = 3 * 3 12 = 3 * 4 to get LCM use any common factors only once so here LCM = 3 * 3 * 4 = 36

OpenStudy (mrm):

The GCD is the greatest common denominator, also known as the greatest common factor GCF. If we look at 9 we see that it is divisible by 1,3, and 9. If we look at 12 we see it is divisible by 1,2,3,4,6, and 12. So we see our GCD is 3. 36 is right but I want you to know why it's right.

OpenStudy (liz_beauty45):

ohhh i see.

OpenStudy (mrm):

So by our formula... \[LCM = \frac{ ab }{ GCD(a,b) } = \frac{ (9)(12) }{ GCD(9,12) } = \frac{ 108 }{ 3 } = 36\]

OpenStudy (mrm):

This always works for any two numbers.

OpenStudy (liz_beauty45):

that makes a lot more since now thanks!

OpenStudy (liz_beauty45):

i just did not know at 1st

OpenStudy (mrm):

That's what learning is all about!

OpenStudy (liz_beauty45):

Thanks again than,ill fan u in a min